OpenStudy (anonymous):

Help me with this, please. Show that in R^n, the lines represented by su1 + u0 and tv1 + v0 are the same if and only if both v1 and v0 - u0 are numberical multiples of u1. (u1, u0, v0, v1 are all vectors, btw).

OpenStudy (helder_edwin):

Let \[ \large L_1=\{s\cdot u_1+v_0:s\in\mathbb{R}\}\quad\text{and}\quad L_2=\{t\cdot v_1+v_0:t\in\mathbb{R}\}. \] First, let's suppose that there exist real numbers \(\alpha,\beta\) such that \[ \large v_1=\alpha u_1\quad\text{and}\quad v_0-u_0=\beta u_1. \]

OpenStudy (helder_edwin):

then, if \(v\in L_2\) we have that for some real \(t\) \[ \large v=t\cdot v_1+v_0=t(\alpha u_1)+\beta u_1+u_0= (t\alpha+\beta)u_1+u_0\in L_1. \] Also, if (\w\in L_1\) we have that for some real \(s\) \[ \large w=s\cdot u_1+u_0=s(1/\alpha\cdot v_1)+v_0-\beta u_1 =(s/\alpha-\beta)+v_0\in L_2 \]. Therefore \(L_1=L_2\).

OpenStudy (helder_edwin):

if have a typo in my first post. it should be \[ \large L_1=\{s\cdot u_1+u_0:s\in\mathbb{R}\}. \]

OpenStudy (helder_edwin):

got it so far?

OpenStudy (anonymous):

Not really, I'm trying to copy that and put in equation but it's not working...

OpenStudy (helder_edwin):

use pencil and paper then

OpenStudy (anonymous):

I got that, thanks.

OpenStudy (helder_edwin):

try and do the reciprolycal by yourself. (it is an if-and-only-if theorem)

OpenStudy (anonymous):

Yes, I think I will.